The admission fee at an amusement park is 1.5 dollars for children and 4 dollars for adults. On a certain day, 304 people entered the park, and the admission fees collected totaled 906 dollars. How many children and how many adults were admitted?

Answer by Boreal(15235)

A+C=304
4A+1.5C=906
multiply first by -4 and eliminate A
-4A-4C=-1216
-2.5C=-310
C=124 for $186
A=180 for $720